OpenStudy (anonymous):

the probability that a professional baseball player will get a hit is 1/3. calculate the exact probability that he will get at least 3 hits in 5 attempts

OpenStudy (amistre64):

mmmmm mmmmh mmmhm mmhmm mhmmm hmmmm mmmhh mmhmh mhmmh hmmmh mmhhm mhmhm hmmhm mhhmm hmhmm hhmmm hhhmm hhmhm hmhhm mhhhm hhmmh hmhmh = 10 (1*1*1*2*2/3^5) mhhmh hmmhh mhmhh mmhhh hhhhm hhhmh hhmhh = 5(1*1*1*1*2/3^5) hmhhh mhhhh hhhhh = 1(1*1*1*1*1/3^5) and there is a total of 32 ways to hit and miss a ball in 5 swings \[\frac{10}{32}\frac{4}{3^5}+\frac{5}{32}\frac{4}{3^5}\] \[\frac{10}{32}\frac{4}{3^5}+\frac{5}{32}\frac{2}{3^5}+\frac{1}{32}\frac{1}{3^5}\]